OpenStudy (anonymous):

can you help please. 3/8w - 22 = -9/8w +4

hero (hero):

Use the drawing feature to write the equations

OpenStudy (anonymous):

\[\frac{ 3 }{ 8}w -22 = -\frac{ 9 }{ 8 }w + 4\]

OpenStudy (precal):

multiply it all by 8 to get rid of the fractions

OpenStudy (anonymous):

ok thank you but what is the answer please , i don't understand

hero (hero):

If you have to ask for the answer, you're definitely not getting it.

OpenStudy (precal):

ok finish solving for the variable

OpenStudy (anonymous):

ok thank you

OpenStudy (precal):

yw

hero (hero):

That way works, but it is a cop out for using the proper method.